# Ryan Martin   and   Jonathan P. Williams

### v1

@@ Abstract
The inferential model (IM) framework offers an alternative to the classical probabilistic (e.g., Bayesian and fiducial) uncertainty quantification in statistical inference. A key distinction is that classical uncertainty quantification takes the form of precise probabilities and offers only limited large-sample validity guarantees, whereas the IM’s uncertainty quantification is imprecise in such a way that exact, finite-sample valid inference is possible. But is the IM’s imprecision and finite-sample validity compatible with statistical efficiency? That is, can IMs be both finite-sample valid and asymptotically efficient? This paper gives an affirmative answer to this question via a new possibilistic Bernstein–von Mises theorem that parallels a fundamental Bayesian result. Among other things, our result shows that the IM solution is efficient in the sense that, asymptotically, its credal set is the smallest that contains the Gaussian distribution with variance equal to the Cramér–Rao lower bound. Moreover, a corresponding version of this new Bernstein–von Mises theorem is presented for problems that involve the elimination of nuisance parameters, which settles an open question concerning the relative efficiency of profiling-based versus extension-based marginalization strategies.
@@ 1 Introduction
Still, the above Bayesian and fiducial developments have not proved to be fully satisfactory, perhaps because large-sample confidence sets fail to tell the whole story. Indeed, the false confidence theorem (Balch et al., 2019; Martin, 2019; Carmichael and Williams, 2018) implies that, whether $n$ is small or large, there exists false hypotheses to which the (Bayesian or fiducial) posterior distribution will tend to assign relatively high probability or confidence. False confidence is related to other familiar-yet-mysterious phenomena such as Stein’s paradox (Stein, 1959), the non-existence of finite-length confidence intervals (Gleser and Hwang, 1987), and that the “confidence” properties enjoyed by confidence distributions are not preserved under probability calculus (e.g., Fraser, 2013, 2011; Schweder and Hjort, 2013). To eschew the inherent unreliability revealed by the false confidence theorem, precise posterior probabilities must be replaced by imprecise probabilities. The inferential model (IM) framework, including the original developments in Martin and Liu, (2013, 2015) and the more recent generalizations in Martin, 2021a ; Martin, 2022a ; Martin, 2022b , does just this. Their proposal is provably valid in a sense that implies safety from false confidence and, e.g., it provides exact finite-sample confidence sets. But one might guess that in order to be provably valid, the IM must sacrifice on efficiency—that is, the aforementioned IM confidence sets must be larger than the asymptotically efficient Bayesian or fiducial credible sets. The examples in the aforementioned references all demonstrate, however, that, despite the imprecision and exact validity, the IM’s solutions are no less (and sometimes more) efficient than their Bayesian and fiducial counterparts. But so far there are no general theoretical results concerning the IM’s efficiency.
@@ 2.1.1 Generalities
So, where ordinary probability is determined by integrating a density function, possibility is determined by optimizing a contour. The values $\{\underline{\Pi}(B),\overline{\Pi}(B)\}$ are often interpreted subjectively, either as (coherent) upper and lower probabilities associated with the proposition “$Z\in B$”—i.e., as maximum buying price and minimum selling price, respectively, for gambles \$$\$1(Z\in B)$ in the sense of Walley, (1991)—or as Shaferian degrees of belief. “Coherence” ensures, among other things, that the possibility measure $\overline{\Pi}$ determines and is determined by its associated credal set (e.g., Levi, 1980, Ch. 4), defined as
@@
The aforementioned limits on the possibility measure’s expressiveness are the following simple constraints: $\underline{\Pi}(B)>0$ implies $\overline{\Pi}(B)=1$ and $\overline{\Pi}(B)<1$ implies $\underline{\Pi}(B)=0$. These limitations do not affect us here, however, because meaningful statistical inferences can be made only when the $\underline{\Pi}$ values are large or the $\overline{\Pi}$ values are small. We will return to this point in Section 2.2 below when we discuss statistical inference.
@@ 2.1.2 Special case: Gaussian possibility
where $F_{D}$ is the ${\sf ChiSq}(D)$ distribution function. The corresponding Gaussian possibility measure, defined via optimization, is denoted by $\overline{\Gamma}_{\mu,\Sigma}$. The notation $\gamma$ and $\overline{\Gamma}$ will be used for the standard Gaussian possibility contour and corresponding possibility measure. A plot of the contour $z\mapsto\gamma(z)$ for $D=1$ can be seen in Figure 1(a) below.
@@ 2.2 Inferential models
The original IM constructions (e.g., Martin and Liu, 2013, 2015) used (nested) random sets and, hence, the connection to possibility theory was indirect. A more streamlined version in Martin, 2022b directly defines the IM’s possibility contour using the probability-to-possibility transform; see, also, Martin, (2015, 2018). An advantage of this new and direct construction is that it avoids the ambiguity in specifying both a data-generating equation and a so-called “predictive random set”—one only needs the model/likelihood function.
@@ 3.3 Main result: possibilistic Bernstein–von Mises
The proof of Theorem 1 is presented in Section 3.8. Before that, however, two technical points are made below. Then the next two subsections offer, first, more context/intuition and, second, a few numerical illustrations of the main result.
As promised, there are equivalent ways to represent the asymptotic approximation of the possibilistic IM’s contour. One such alternative happens to be the most convenient in certain aspects of the proof in Section 3.8 below. Define
@@ 3.6 Implications for uncertainty quantification
Under the conditions of Theorem 1, let $H$ be a fixed hypothesis with $\Theta$ in its interior. Then $\overline{\Pi}_{X^{n}}(H)\to 1$ and $\overline{\Pi}_{X^{n}}(H^{c})\to 0$ in $\mathsf{P}_{\Theta}$-probability.
@@
Under the conditions of Theorem 1, let $f$ be a continuous function on a domain that contains $\Theta$ in its interior. Then $\overline{\Pi}_{X^{n}}f\to f(\Theta)$ in $\mathsf{P}_{\Theta}$-probability.
@@ 3.7 Technical remarks
3.8 Proof of Theorem 1
@@
To kick things off, fix $z$ and set $\theta_{n}^{z}=\Theta+n^{-1/2}z$. Then: This decomposition suggests a two-step proof: Lemmas 1 and 2 below show that the first and second terms are suitably vanishing, which will prove the claim.
@@
4 Cases involving nuisance parameters
4.1 Setup
[…]
4.2 Eliminating nuisance parameters
4.2.1 Extension
4.2.2 Profiling
Recall that the relative likelihood for $\Theta$ is
The profile relative likelihood for $\Phi=f(\Theta)$ is given by
4.2.3 Quick illustration
(a) Joint contour for $(\Phi,\Lambda)$
(b) Marginal contour for $\Phi$
4.2.4 Other strategies
4.3 Background
This is often called the efficient Fisher information matrix. Notice that this matrix is generally smaller (relative to Lowener order) than the matrix $I_{11}(\Phi,\Lambda)$; this difference in the two information measures is the penalty resulting from $\Lambda$ being unknown rather than known, as the latter information matrix unrealistically assumes. Note that the two matrices agree, i.e., $\tilde{I}_{\Phi,\Lambda}=I_{11}(\Phi,\Lambda)$, when $I_{12}(\Phi,\Lambda)=0$; this corresponds to what is commonly referred to as parameter orthogonality, which we discuss this further following the statement (and proof sketch) of Theorem 2 below.
4.4 Marginal possibilistic Bernstein–von Mises
Theorem 2 (Profiling-based marginalization).
[…]
for any compact $K\subset\mathbb{R}^{D_{\phi}}$.
@@ 4.4 Marginal possibilistic Bernstein–von Mises
Theorem 3 (Extension-based marginalization).
[…]
for any compact $K\subset\mathbb{R}^{D_{\phi}}$.
The proof of this result is pretty straightforward, at least intuitively. The only thing that $\pi_{X^{n}}^{\text{\sc ex}}(\phi)$ could possibly merge with is the supremum of the limiting version of the original contour, i.e., $\gamma_{X^{n}}(\phi,\lambda)$, over the nuisance parameter $\lambda$. The uniformity established in Theorem 1 makes this argument rigorous. Now it is just a matter of carrying out the optimization in the limiting Gaussian contour $\gamma_{X^{n}}$. But we did this already in Section 2.1—see Equation (2)—when discussing the extension principle applied to Gaussian possibility measures. Indeed, that same argument applied to the particular Gaussian here gives a contour of the form (31), which proves the claim.
4.5 Numerical illustrations
Example 5.
$(X\mid U)\sim{\sf N}_{n}(U\Lambda,\Phi\,I_{n\times n}),$
and, therefore, the relative profile likelihood is
Example 6.
Example 7.
5 Conclusion
Acknowledgments

### v2

@@ Abstract
The inferential model (IM) framework offers an alternative to the classical probabilistic (e.g., Bayesian and fiducial) uncertainty quantification in statistical inference. A key distinction is that classical uncertainty quantification takes the form of precise probabilities and offers only limited large-sample validity guarantees, whereas the IM’s uncertainty quantification is imprecise in such a way that exact, finite-sample valid inference is possible. But are the IM’s imprecision and finite-sample validity compatible with statistical efficiency? That is, can IMs be both finite-sample valid and asymptotically efficient? This paper gives an affirmative answer to this question via a new possibilistic Bernstein–von Mises theorem that parallels a fundamental Bayesian result. Among other things, our result shows that the IM solution is efficient in the sense that, asymptotically, its credal set is the smallest that contains the Gaussian distribution with variance equal to the Cramér–Rao lower bound. Moreover, a corresponding version of this new Bernstein–von Mises theorem is presented for problems that involve the elimination of nuisance parameters, which settles an open question concerning the relative efficiency of profiling-based versus extension-based marginalization strategies.
@@ 1 Introduction
Still, the above Bayesian and fiducial developments have not proved to be fully satisfactory, perhaps because large-sample confidence sets fail to tell the whole story. Indeed, the false confidence theorem (Balch et al., 2019; Martin, 2019; Carmichael and Williams, 2018) implies that, whether $n$ is small or large, there exists false hypotheses to which the (Bayesian or fiducial) posterior distribution will tend to assign relatively high probability or confidence. False confidence is related to other familiar-yet-mysterious phenomena such as Stein’s paradox (Stein, 1959), the non-existence of finite-length confidence intervals (Gleser and Hwang, 1987), and that the “confidence” properties enjoyed by confidence distributions are not preserved under probability calculus (e.g., Fraser, 2013, 2011; Schweder and Hjort, 2013). To eschew the inherent unreliability revealed by the false confidence theorem, precise posterior probabilities must be replaced by imprecise probabilities. The inferential model (IM) framework, including the original developments in Martin and Liu, (2013); Martin and Liu, 2015a and the more recent generalizations in Martin, 2021a ; Martin, 2022a ; Martin, 2022b , does just this. Their proposal is provably valid in a sense that implies safety from false confidence and, e.g., it provides exact finite-sample confidence sets. But one might guess that in order to be provably valid, the IM must sacrifice on efficiency—that is, the aforementioned IM confidence sets must be larger than the asymptotically efficient Bayesian or fiducial credible sets. The examples in the aforementioned references all demonstrate, however, that, despite the imprecision and exact validity, the IM’s solutions are no less (and sometimes more) efficient than their Bayesian and fiducial counterparts. But so far there are no general theoretical results concerning the IM’s efficiency.
@@ 2.1.1 Generalities
So, where ordinary probability is determined by integrating a density function, possibility is determined by optimizing a contour. The values $\{\underline{\Pi}(B),\overline{\Pi}(B)\}$ are often interpreted subjectively, either as (coherent) upper and lower probabilities associated with the proposition “$Z\in B$”—i.e., as maximum buying price and minimum selling price, respectively, for gambles \$$\$1(Z\in B)$ in the sense of Walley, (1991)—or as Shaferian degrees of belief. “Coherence” ensures, among other things, that the possibility measure $\overline{\Pi}$ determines and is determined by its associated credal set (e.g., Levi, 1980, Ch. 4), which is the set of all ordinary/precise probabilities dominated by $\overline{\Pi}$ in the sense that the numerical probabilities they assign to events are no larger than the possibilities assigned to the same events by $\overline{\Pi}$. Mathematically, the credal set is given by
@@
The aforementioned limits on the possibility measure’s expressiveness are the following simple constraints: $\underline{\Pi}(B)>0$ implies $\overline{\Pi}(B)=1$ and $\overline{\Pi}(B)<1$ implies $\underline{\Pi}(B)=0$. This is intuitively clear from the interpretation of $(\underline{\Pi},\overline{\Pi})$ as necessity and possibility: if there is any genuine shred of evidence that implies $B$, then $B$ must be absolutely possible; and if there is any doubt that $B$ is possible, then there can be no genuine shred of evidence that implies $B$. These limitations do not affect us here, however, because meaningful statistical inferences can be made only when the $\underline{\Pi}$ values are large or the $\overline{\Pi}$ values are small. We will return to this point in Section 2.2 below when we discuss statistical inference.
@@ 2.1.2 Special case: Gaussian possibility
where $F_{D}$ is the ${\sf ChiSq}(D)$ distribution function. Note that what we are referring to here as the “Gaussian possibility contour” is different from the Gaussian possibility distribution, with $\pi(z)=\exp\{-\frac{1}{2}(z-\mu)^{\top}\Sigma^{-1}(z-\mu)\}$, that is often found in the literature (e.g., Denœux, 2023b ). The corresponding Gaussian possibility measure, defined via optimization, is denoted by $\overline{\Gamma}_{\mu,\Sigma}$. The notation $\gamma$ and $\overline{\Gamma}$ will be used for the standard Gaussian possibility contour and corresponding possibility measure. A plot of the contour $z\mapsto\gamma(z)$ for $D=1$ can be seen in Figure 1(a) below.
@@ 2.2 Inferential models
The original IM constructions (e.g., Martin and Liu, 2013; Martin and Liu, 2015a ) used (nested) random sets and, hence, the connection to possibility theory was indirect. A more streamlined version in Martin, 2022b directly defines the IM’s possibility contour using the probability-to-possibility transform; see, also, Martin, (2015, 2018). An advantage of this new and direct construction is that it avoids the ambiguity in specifying both a data-generating equation and a so-called “predictive random set”—one only needs the model/likelihood function.
@@ 3.3 Main result: possibilistic Bernstein–von Mises
The proof of Theorem 1 is presented in Appendix A.1. Before that, however, two technical points are made below. Then the next two subsections offer, first, more context/intuition and, second, a few numerical illustrations of the main result.
As promised, there are equivalent ways to represent the asymptotic approximation of the possibilistic IM’s contour. One such alternative happens to be the most convenient in certain aspects of the proof in Appendix A.1 below. Define
@@ 3.6 Implications for uncertainty quantification
Let $H$ be a fixed hypothesis with $\Theta$ in its interior. Under the conditions of Theorem 1, $\overline{\Pi}_{X^{n}}(H)\to 1$ and $\overline{\Pi}_{X^{n}}(H^{c})\to 0$ in $\mathsf{P}_{\Theta}$-probability.
@@
Let $f$ be a continuous function on a domain that contains $\Theta$ in its interior. Under the conditions of Theorem 1, $\overline{\Pi}_{X^{n}}f\to f(\Theta)$ in $\mathsf{P}_{\Theta}$-probability.
@@ 3.7 Technical remarks
4 Cases involving nuisance parameters
4.1 Setup
[…]
4.2 Eliminating nuisance parameters
4.2.1 Extension
4.2.2 Profiling
Recall that the relative likelihood for $\Theta$ is
The profile relative likelihood for $\Phi=f(\Theta)$ is given by
4.2.3 Quick illustration
(a) Joint contour for $(\Phi,\Lambda)$
(b) Marginal contour for $\Phi$
4.2.4 Other strategies
In a different but related direction, the approach of Martin and Liu, 2015b offers a novel perspective on the elimination of nuisance parameters. The authors’ proposal is quite powerful in that it affords users the flexibility to do more-or-less whatever it takes—e.g., model-based strategies like marginalization and algebraic manipulations of data-generating equations—to eliminate the nuisance parameters and reduce the overall dimension of the problem so that efficient inference is within reach. A disadvantage to their proposal is that, once the nuisance parameter elimination is done, inference requires specification of a suitable “predictive random set.” At present, guidelines on how to choose this random set to achieve both statistical and computational efficiency are lacking.
4.3 Background
This is often called the efficient Fisher information matrix. Notice that this matrix is generally smaller (relative to Loewner order) than the matrix $I_{11}(\Phi,\Lambda)$; this difference in the two information measures is the penalty resulting from $\Lambda$ being unknown rather than known, as the latter information matrix unrealistically assumes. Note that the two matrices agree, i.e., $\tilde{I}_{\Phi,\Lambda}=I_{11}(\Phi,\Lambda)$, when $I_{12}(\Phi,\Lambda)=0$; this corresponds to what is commonly referred to as parameter orthogonality, which we discuss this further following the statement (and proof sketch) of Theorem 2 below.
4.4 Marginal possibilistic Bernstein–von Mises
Theorem 2 (Profiling-based marginalization).
[…]
for any compact $K\subset\mathbb{R}^{D_{\phi}}$.
Proof.
A sketch of the proof is given in Appendix A.2. ∎
Theorem 3 (Extension-based marginalization).
[…]
for any compact $K\subset\mathbb{R}^{D_{\phi}}$.
Proof.
See Appendix A.3. ∎
4.5 Numerical illustrations
Example 5.
$(X\mid U)\sim{\sf N}_{n}(U\Lambda,\Phi\,I_{n\times n}),$
and, therefore, the relative profile likelihood is
Example 6.
Example 7.
5 Conclusion
Acknowledgments
Appendix A Technical details
A.1 Proof of Theorem 1
@@
This decomposition suggests a two-step proof: Lemmas 1 and 2 below show that the first and second terms are suitably vanishing, which will prove the claim.
@@
A.2 Proof of Theorem 2: sketch
@@
A.3 Proof of Theorem 3
The proof is pretty straightforward. The only thing that $\pi_{X^{n}}^{\text{\sc ex}}(\phi)$ could possibly merge with is the supremum of the limiting version of the original contour, i.e., $\gamma_{X^{n}}(\phi,\lambda)$, over the nuisance parameter $\lambda$. The uniformity established in Theorem 1 makes this argument rigorous. Now it is just a matter of carrying out the optimization in the limiting Gaussian contour $\gamma_{X^{n}}$. But we did this already in Section 2.1—see Equation (2)—when discussing the extension principle applied to Gaussian possibility measures. Indeed, that same argument applied to the particular Gaussian here gives a contour of the form (27), which proves the claim.